# Planning of Heuristics: Strategic Planning on Large Language Models with Monte Carlo Tree Search for Automating Heuristic Optimization

### v1

@@ 3.2 PoH Framework
PoH regards the heuristic optimization problem as a Markov Decision Process (MDP), defined by the tuple  $(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R})$, with a defined state and action space. Here, $\mathcal{S}$ represents the state space, $\mathcal{A}$ the action space, $\mathcal{T}$ the transition function $\mathcal{T}:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{S}$, and $\mathcal{R}$ the reward function $\mathcal{R}:\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}$. As illustrated in Figure 1 (a), given a current state $s_{t}$, PoH iteratively generates an action $a_{t}$ according to $a_{t}\sim p_{\mathcal{B}}(a|s_{t})$. The action generation process, detailed in Figure 1 (b), comprises two steps: first, evaluating the heuristic (state) generated by the base model. This evaluation involves replacing the distance matrix update function of the GLS search algorithm with the generated heuristic and then evaluating it on the training set to obtain the current heuristic’s optimal gap (%). Subsequently, for action generation, the LLM optimizer generates improvement suggestions based on the current heuristic and the optimal gap (%). PoH then determines the subsequent state using the transition function $p_{\mathcal{O}}(s_{t+1}|s_{t},a_{t})$ to update the heuristic. Given the current improvement suggestions (action), the LLM optimizer $p_{\mathcal{O}}$ generates a new heuristic (state) that incorporates relevant domain knowledge and effectively addresses model suggestions, which is similar to how AHD via Hyper-Heuristics revises algorithms based on improvement suggestions.
The quality of the resulting state $s_{t}$ after applying the action $a_{t}$ is then evaluated by the reward function $r_{t}=r(s_{t},a_{t})$. The reward function is defined as 1 minus the percentage gap between the solution and the optimal solution for an instance, aiming to maximize solution quality and minimize the gap with the optimal solution.
@@ 3.3 Planning with Monte Carlo Tree Search
Selection Starting from the root node, the selection phase traverses the child nodes according to the Upper Confidence Bound (UCT) formula, which balances exploitation and exploration, until a leaf node is reached. The UCT formula is as follows:
@@
Expansion When a leaf node is reached and a terminal state has not yet been achieved, the expansion phase creates one or more new nodes. To derive diverse improvement suggestions (actions), we may sample multiple training batches. Among these new nodes, the one with the highest reward is then passed to the next simulation step.
Simulation Starting from the current node, the simulation phase selects actions from all possible actions according to a policy, resulting in state transitions. If a terminal state is not reached, actions continue to be selected until a terminal state is reached. To improve efficiency in our experiments, we select the action with the highest local reward from the multiple actions generated during the expansion phase.
Back-propagation Back-propagation propagates the simulation result back through the search tree to update node information. When the simulation reaches a terminal state (e.g., reaching maximum depth or satisfying a predefined termination criterion), the $Q$ value of the current node is updated. By continuously updating node information, MCTS guides the search process in more promising directions.
PoH iteratively executes these four phases to explore the heuristic space. The algorithm terminates when a predefined number of iterations is reached, and the best trajectory and corresponding node are then selected for evaluation.

### v2

@@ 3.2 PoH Framework
PoH regards the heuristic optimization problem as a Markov Decision Process (MDP), defined by the tuple  $(\mathcal{S},\mathcal{A},\mathcal{T},\mathcal{R})$, with a defined state and action space. Here, $\mathcal{S}$ represents the state space, $\mathcal{A}$ is the action space, $\mathcal{T}$ denotes the transition function, $\mathcal{T}:\mathcal{S}\times\mathcal{A}\mapsto\mathcal{S}$, and $\mathcal{R}$ is the reward function $\mathcal{R}:\mathcal{S}\times\mathcal{A}\mapsto\mathbb{R}$. In our work, we formulate the MDP concretely as follows.
State $\mathcal{S}$ is a set of generated heuristics, where $s_{t}\in\mathcal{S}$, and $s_{t}$ is the generated heuristic at step $t$.
Action $\mathcal{A}$ is a set of improvement suggestions, where $a_{t}\in\mathcal{A}$ and $a_{t}$ is the generated improvement suggestion for heuristic $s_{t}$.
Transition function $\mathcal{T}$: When applying an action $a_{t}$ to state $s_{t}$, it yields $s_{t+1}$, which means the heuristic (state $s_{t}$) accepts the improvement suggestion (action $a_{t}$) and updates to a new heuristic (state $s_{t+1}$).
Reward $\mathcal{R}$ is evaluated as $1-\frac{O_{st}-O_{best}}{O_{best}}$ , where $O_{best}$ is known instance-specific baselines, $O_{st}$ is the current heuristic.
As illustrated in Figure 1 (a), given a current state $s_{t}$, PoH iteratively generates an action $a_{t}$ according to $a_{t}\sim p_{\theta}(a|s_{t})$, $p_{{B}}(a|s_{t})$ represents that the base LLM $p_{\theta}$ generates improvement actions $a$ given the current state $s_{t}$. The action generation process, detailed in Figure 1 (b), comprises two steps: first, evaluating the heuristic (state) generated by the base model. This evaluation involves replacing the distance matrix update function of the GLS search algorithm with the generated heuristic and then evaluating it on the training set to obtain the current heuristic’s optimal gap (%). Subsequently, for action generation, the optimizer LLM $p_{\mathcal{B}}$ generates improvement suggestions $a_{t}$ based on the current heuristic $s_{t}$ and the optimal gap (%). PoH then determines the subsequent state using the transition function $p_{\mathcal{\theta}}(s_{t+1}|s_{t},a_{t})$ to update the heuristic. Given the current improvement suggestions (action), the optimizer LLM $p_{\mathcal{\theta}}$ generates a new heuristic (state) that incorporates relevant domain knowledge and effectively addresses model suggestions, which is similar to how AHD via Hyper-Heuristics updates algorithms based on improvement suggestions.
The quality of applying the action $a_{t}$ to state $s_{t}$ is then evaluated by the reward function $r_{t}=r(s_{t},a_{t})$. The reward function is defined as 1 minus the percentage gap between the solution and the optimal solution for an instance, aiming to maximize solution quality and minimize the gap with the optimal solution.
@@ 3.3 Planning with Monte Carlo Tree Search
PoH incorporates MCTS as its strategic planning method. A typical MCTS loop consists of the following 4 steps.
Selection Starting from the root node in a search tree, the selection phase traverses the child nodes according to the Upper Confidence Bound (UCT) formula, which balances exploitation and exploration, until a leaf node is reached. The UCT formula is as follows:
@@
Expansion When a leaf node is reached and a terminal state has not yet been achieved, the expansion phase creates one or more new nodes. These new nodes represent future states that may be reached from the current state. During the expansion phase, no evaluation or simulation of these newly added child nodes takes place. Expansion merely increases the structure of the tree in preparation for subsequent simulation steps.
Simulation Starting from the current node, the simulation phase selects actions from all possible actions according to a policy, resulting in state transitions. If a terminal state is not reached, actions continue to be selected until a terminal state is reached.
Back-propagation Back-propagation propagates the simulation result back through the search tree to update node information. When the simulation reaches a terminal state, the $Q$ value of the current node is updated. By continuously updating node information, MCTS guides the search process in more promising directions.
$Q(s_{t},a_{t})=\max_{s_{t},a_{t},r_{t},\dots,s_{l},a_{l},r_{l},s_{l+1}}\operatorname{avg}(r_{t},\dots,r_{l}).$ (2)
Above all, the pseudo code of the PoH method is given as Algorithm 1. The terminal state is when it reaches the predefined maximum depth or meets the early-stopping condition (Line 18 in Algorithm 1). The early stopping condition is activated when the state’s reward is below a minimum threshold or above a maximum threshold. Specifically, the minimum threshold is the average of the rewards of the parent node and the root node, while the maximum threshold is the maximum value among all current nodes. PoH iteratively executes these four phases to explore the heuristic space. The algorithm terminates when a predefined number of iterations is reached, and the best trajectory and corresponding node are then selected for evaluation.